# The minimal periodicity for integral bases of pure number fields

### v1

@@ Abstract.
Fix $n\geq 3$. For the pure field $K_{a}=\mathbb{Q}(\theta)$, $\theta^{n}=a$ with $a$ $n$th-power-free, we encode an integral basis in the fixed coordinate $\{1,\theta,\dots,\theta^{n-1}\}$ by its shape. We prove a sharp local-to-global principle: for each $p^{e}\!\parallel n$, the local shape at $p$ is determined by $a\bmod p^{\,e+1}$, and this precision is optimal. Moreover, the global shape is periodic with minimal modulus
@@ 1. Introduction
A classical problem in algebraic number theory is to study the ring of integers $\mathcal{O}_{K}$ of a number field $K$, in particular for the pure fields $K=\mathbb{Q}(\theta)$ with $\theta^{n}=a$ (see [Gaá19] for a modern treatment). Recently, two complementary phenomena have guided works on this direction. First, in small degrees ($3\leq n\leq 9$) Gaál–Remete [GR17] showed that the global pattern of integral bases is periodic as $a$ varies, with period $n^{2}$. In particular, if one writes an integral basis with rational coefficients in the $\{1,\theta,\dots,\theta^{n-1}\}$-coordinates, then every prime dividing a denominator must divide $n$ (a global restriction that already foreshadows a local explanation). This offers a striking, hands‑on picture of regularity in families of pure fields. Secondly, for general $n$ and squarefree $a$, Jakhar–Khanduja–Sangwan [JKS21] gave explicit integral bases and proved a global period $n\cdot\prod_{p\mid n}p$. Their construction is intrinsically local: it splices together $p$-integral bases at primes $p\mid n$ and encodes, for each $m=1,\dots,n-1$, a pp‑adic denominator exponent and a small “linear correction”. The dependence on $a$ is controlled by a short list of pp‑adic invariants (e.g. $v_{p}(a^{p-1}-1)$), revealing that only one extra pp‑adic digit beyond $p^{e_{p}}\parallel n$ is ever needed to fix the local shape at $p$.
In this paper we answer the question on how much information about $a$ do the local computations at each $p\mid n$ really need? Equivalently, what is the smallest modulus $M(n)$ such that, whenever $a\equiv a^{\prime}\pmod{M(n)}$, the structure of the integral basis of $\mathcal{O}_{\mathbb{Q}(\sqrt[n]{a})}$ is the same as for $a^{\prime}$? We show that the Jakhar–Khanduja–Sangwan period is in fact minimal, in a precise sense, for all $n\geq 3$, and we do so in a way that emphasizes the underlying local mechanism.mThe first novelty of our paper is we unify and formalize these two phenomena by introducing the concept of shape (defined precisely in  2.2) as the sequence, for $m=1,\dots,n-1$, consisting of (i) the prime‐power denominators supported on $p\mid n$ and (ii) the corresponding linear corrections recorded modulo those denominators. We then prove that (see Theorem 3.1) the smallest global modulus governing periodicity of the shape is always
@@
Our proof is short and local-to-global. More precisely, fix a prime power $p^{e}\parallel n$. Running the Ore–Montes (or higher Newton polygon) procedure (see [MN92, GMN12, GMN15]) for $f(x)=x^{n}-a$ at $p$ up to order $e$ shows that the entire $p$-type (slopes, residual polynomials, and hence the $p$-integral basis with its denominators/corrections) is only determined by the residue class
$a\pmod{p^{\,e+1}}$.
Already among units $a\in\mathbb{Z}_{p}^{\times}$ one can choose
@@
for which the numerical invariant $r_{p}=v_{p}(a^{p-1}-1)-1$ jumps from e1e-1 to $\geq e$. Since the local denominators and linear corrections depend on $d_{p}=\min\{r_{p},e\}$ [JKS21, Theorem 1.6], the $p$-part of the shape changes; hence $p^{e+1}$ is the least local modulus at $p$.
By the Chinese Remainder Theorem the global shape depends only on $a\bmod\prod_{p\mid n}p^{e_{p}+1}=n\cdot\mathrm{rad}(n)$, giving the optimal upper bound. The local sharpness at every $p\mid n$ makes this modulus minimal.
@@
The point of view developed here—the local shape at each $p\mid n$ and its global gluing—yields several immediate applications. Firstly, we can show that one can compute, for any fixed $n$, a finite lookup table indexed by the classes $a\bmod M(n)$ recording the pp–denominator exponents and reduced corrections for every $p\mid n$. Thanks to Theorem 3.1, the table itself can be filled entirely locally by running the higher Newton polygon up to order $e_{p}$ at each $p\mid n$. This offers a uniform and predictable alternative to ad hoc casework used in small degrees.
The second application concerns monogeneity tests localized at the primes $p\mid n$. When $p\parallel n$, the obstruction coming from the order‑$1$ Newton polygon at $p$ is governed by the single “Wieferich” congruence $a^{p-1}\equiv 1\pmod{p^{2}}$: fixing $a\bmod p$ with $a^{p-1}\equiv 1\bmod p$, the dichotomy between $a^{p-1}\not\equiv 1\pmod{p^{2}}$ and $a^{p-1}\equiv 1\pmod{p^{2}}$ decides whether the order‑$1$ type is pp‑regular and hence whether the local index contribution at $p$ vanishes. More generally, for $p^{e}\parallel n$ and $v_{p}(a)=0$ the relevant invariant is $d_{p}(a)=\min\{\,v_{p}(a^{p-1}-1)-1,\,e\,\}$, and Theorem 3.1 shows that this (and with it all pp–denominator exponents and reduced corrections in an explicit basis) is determined by $a\bmod p^{e+1}$ and cannot be recovered from $a\bmod p^{e}$ alone.
@@
1.1. Historical remarks and related results
@@
When Dedekind’s criterion fails (i.e. some $p\mid[\mathcal{O}_{K}:\mathbb{Z}[\theta]]$), the modern local machine is based on Newton polygons at $p$, initiated by Ore in the 1920s and refined into a systematic algorithm by Montes and collaborators: one builds (higher-order) $p$-adic Newton polygons, reads residual polynomials along principal sides, and from this recovers $p$-adic factorization, indices, discriminants, and $p$-integral bases [Ore28, MN92, GMN12, GMN15]. This underlies much current algorithmic number theory around integral bases and prime decomposition. For practical $p$-integral basis extraction and index computations using Newton polygons, see ElFadil and his collaborators’ papers [EFMN12, EF20] and expositions cited therein. For the complexity and implementation discussions, take a look at [BNS13]. These provide worked implementations that agree with and conceptually support the p-local pieces used by Gaál–Remete and Jakhar–Khanduja–Sangwan.
@@
Section 2 introduces the shape of an integral basis and gives illustrative examples. Section 3 is preceded by our main local–to–global result (Theorem 3.1) and its direct consequences. We also include a tame relative version (Proposition 3.6) and examples. Section 4 develops applications: sharp local monogeneity tests (including the Wieferich threshold at $p\parallel n$), explicit discriminant jumps, and counting/density results for prescribed shape classes, with worked quartic and sextic cases. An Appendix A summarizes the Ore–Montes toolkit specialized to $x^{n}-a$ for quick reference.
@@ Acknowledgments
I am grateful to the Morningside Center of Mathematics, Chinese Academy of Sciences, for its support and a stimulating research environment.
@@ 2. Shape, periodicity, and minimal period
Throughout let $n\geq 3$ and $K_{a}=\mathbb{Q}(\theta)$ with $\theta^{n}=a$, where $a\in\mathbb{Z}$ is $n$th‑power‑free. Write $n=\prod_{p}p^{e_{p}}$ and $\mathrm{rad}(n)=\prod_{p\mid n}p$. When we write an integral basis in the $\{1,\theta,\dots,\theta^{n-1}\}$–coordinates we mean a $\mathbb{Z}$–basis of the form
@@
with $D_{m}\in\mathbb{Z}_{\geq 1}$, $\beta_{m}\in\mathbb{Z}[\theta]$ and linear in $\{1,\theta,\dots,\theta^{m-1}\}$. We use the following standard assumption in [JKS21] under which the explicit integral bases of $\mathcal{O}_{K_{a}}$ are stated. In particular, it holds whenever $a$ is squarefree or $(a,n)=1$.
@@
Write $n=\prod_{p}p^{e_{p}}$ and the factorization $a=\prod_{j=1}^{n-1}a_{j}$ with the $a_{j}$ squarefree and pairwise coprime. For $0\leq m\leq n-1$ set
@@
For each $m$ and each $p\mid n$, let $k_{p,m}$ be the largest integer $k$ with $0\leq k\leq d_{p}(a)$ such that $m\geq n-n/p^{\,k}$ (put $k_{p,m}=0$ if no such $k$ exists). We obtain the following remarkable result.
@@
If $k_{p,m}=0$ we regard the residue class $\beta_{m}\bmod p^{\,k_{p,m}}$ as trivial; in particular the pp–local shape depends only on those $\beta_{m}$ with $k_{p,m}>0$.
@@
We say the shape is periodic modulo $N$ if $\mathbf{S}(a+N)=\mathbf{S}(a)$ for all $n$th–power–free $a\in\mathbb{Z}$. A positive integer $N$ with this property is a period (for fixed $n$).
@@
Assume Hypothesis (H). Write $r=v_{2}(a-1)-1$ when $a$ is odd and r=1r=-1 when $2\mid a$; put $d=\min\{r,2\}$.
Let $a$ be 4th‑power‑free. Put $r=v_{2}(a-1)-1$ when $a$ is odd, and r=1r=-1 when $2\mid a$. Then $d=\min\{r,2\}$. The general rule above gives
@@ 3. Main theorems
The main tool we use is the Ore–Montes machinery, referring to Guardia–Montes–Nart [GMN12]. Fix $p\mid n$ with $p^{e}\parallel n$ and write $v=v_{p}$. Consider $f(x)=x^{n}-a\in\mathbb{Z}[x]$. Given a monic $\varphi\in\mathbb{Z}_{p}[x]$ with irreducible reduction, every $f$ has a unique $\varphi$‑adic expansion
$f(x)=\sum_{i=0}^{\ell}a_{i}(x)\,\varphi(x)^{i},\quad a_{i}\in\mathbb{Z}_{p}[x],\ \deg a_{i}<\deg\varphi,$
whose principal $\varphi$‑Newton polygon $N^{-}_{\varphi}(f)$ is the lower convex hull of the points $(i,u_{i})$ with $u_{i}:=v(a_{i})$ (points with $u_{i}=\infty$ ignored). The residual coefficients along a side $S$ of slope $\lambda=-h/e$ are defined by
$c_{s+je}\ :=\ \overline{\,a_{s+je}/p^{\,u_{s}-jh}\,}\ \in\ \mathbb{F}_{\varphi},\qquad j=0,\dots,d,$
and the associated residual polynomial is $R_{\lambda}(f)(y)=\sum_{j=0}^{d}c_{s+je}y^{j}\in\mathbb{F}_{\varphi}[y]$ (here $(s,u_{s})$ is the initial point of $S$ and $d$ its degree).
Higher orders are built by iterating this construction with types
$t=(\varphi_{1};\lambda_{1},\varphi_{2};\ldots;\lambda_{r-1},\varphi_{r};\psi_{r}).$
At order $r$ one considers the $\varphi_{r}$–expansion with respect to the augmented valuation $\omega_{r}$ determined by $t$, forms the principal polygon $N^{-}_{r}(f)$ and residual polynomials $R_{r}(f)$, and so on. The key facts are as follows (see [GMN12, §2-4]):
  1. (1)
    the order–$r$ residual coefficients are again reductions of certain pp–adic coefficients divided by a power of $p$ determined by the initial ordinate of the relevant side;
  2. (2)
    once the (finite) chain of polygons and residual polynomials is fixed up to the necessary order, the pp–factorization of $p\mathcal{O}_{K_{a}}$ and a pp–integral basis follow from the index theorem and the pp–integral–basis construction.
@@
Let $n\geq 3$, $a\in\mathbb{Z}$ be $n$th–power–free, and $K_{a}=\mathbb{Q}(\theta)$ with $\theta^{n}=a$ satisfying hypothesis (H). For each prime $p\mid n$ write $p^{e_{p}}\parallel n$ and let $\mathbf{S}_{p}(a)$ and $\mathbf{S}(a)$ be the local and global shapes.
@@
    (Local determinacy.) For every $p\mid n$, the local shape $\mathbf{S}_{p}(a)$ depends only on the residue class of $a$ modulo $p^{\,e_{p}+1}$. Equivalently, if $a\equiv a^{\prime}\pmod{p^{\,e_{p}+1}}$, then $\mathbf{S}_{p}(a)=\mathbf{S}_{p}(a^{\prime})$.
@@
    and $M(n)$ is the least such modulus: if $N\geq 1$ has the property that $a\equiv a^{\prime}\pmod{N}$ implies $\mathbf{S}(a)=\mathbf{S}(a^{\prime})$ for all $n$th–power–free $a,a^{\prime}$, then $M(n)\mid N$.

### v2

@@ Abstract.
Fix $n\geq 3$. For the pure field $K_{a}=\mathbb{Q}(\theta)$ with $\theta^{n}=a$, where $a\neq\pm 1$ is $n$th-power-free, we encode an integral basis in the fixed coordinate $\{1,\theta,\dots,\theta^{n-1}\}$ by its shape. We prove a sharp local-to-global principle: for each $p^{e}\!\parallel n$, the local shape at $p$ is determined by $a\bmod p^{\,e+1}$, and this precision is optimal. Moreover, the global shape is periodic with minimal modulus
@@ 1. Introduction
A classical problem in algebraic number theory is to study the ring of integers $\mathcal{O}_{K}$ of a number field $K$, in particular for the pure fields $K=\mathbb{Q}(\theta)$ with $\theta^{n}=a$ (see [GAÁ19] for a modern treatment). Throughout we assume that $a$ is $n$th-power-free, $a\neq\pm 1$, and that $x^{n}-a$ is irreducible over $\mathbb{Q}$ (equivalently, $K_{a}:\mathbb{Q}]=n$). Recently, two complementary phenomena have guided works on this direction. First, in small degrees ($3\leq n\leq 9$) Gaál–Remete [GR17] showed that the global pattern of integral bases is periodic as $a$ varies, with period $n^{2}$. In particular, if one writes an integral basis with rational coefficients in the $\{1,\theta,\dots,\theta^{n-1}\}$-coordinates, then every prime dividing a denominator must divide $n$ (a global restriction that already foreshadows a local explanation). This offers a striking, hands‑on picture of regularity in families of pure fields. Secondly, for general $n$ and squarefree $a$, Jakhar–Khanduja–Sangwan [JKS21] gave explicit integral bases and proved a global period $n\cdot\prod_{p\mid n}p$. Their construction is intrinsically local: it splices together $p$-integral bases at primes $p\mid n$ and encodes, for each $m=1,\dots,n-1$, a pp‑adic denominator exponent and a small “linear correction”. The dependence on $a$ is controlled by a short list of pp‑adic invariants (e.g. $v_{p}(a^{p-1}-1)$), revealing that only one extra pp‑adic digit beyond $p^{e_{p}}\parallel n$ is ever needed to fix the local shape at $p$.
In this paper we answer the question on how much information about $a$ do the local computations at each $p\mid n$ really need? Equivalently, what is the smallest modulus $M(n)$ such that, whenever $a\equiv a^{\prime}\pmod{M(n)}$, the structure of the integral basis of $\mathcal{O}_{\mathbb{Q}(\sqrt[n]{a})}$ is the same as for $a^{\prime}$? We show that the Jakhar–Khanduja–Sangwan period is in fact minimal, in a precise sense, for all $n\geq 3$, and we do so in a way that emphasizes the underlying local mechanism. The first novelty of our paper is we unify and formalize these two phenomena by introducing the concept of shape (defined precisely in  2.2) as the sequence, for $m=1,\dots,n-1$, consisting of (i) the prime‐power denominators supported on $p\mid n$ and (ii) the corresponding linear corrections recorded modulo those denominators. We then prove that (see Theorem 3.1) the smallest global modulus governing periodicity of the shape is always
@@
Our proof is short and local-to-global and uses only the explicit $p$-local invariants and bases of Jakhar–Khanduja–Sangwan [JKS21] under Hypothesis (H). Fix a prime power $p^{e}\parallel n$ and write $v_{p}$ for the $p$-adic valuation. When $v_{p}(a)=0$ we set
$r_{p}(a)=v_{p}(a^{p-1}-1)-1,\qquad d_{p}(a)=\min\{r_{p}(a),e\},$
and when $v_{p}(a)>0$ we have $r_{p}(a)=-1$ (hence $d_{p}(a)=-1$). In [JKS21, Theorem 1.6], the pp–denominator exponents $k_{p,m}$ are explicit functions of $p,n,m$ and the truncated invariant $d_{p}(a)=\min\{v_{p}(a^{p-1}-1)-1,e\}$ (in the unit branch $v_{p}(a)=0$). Moreover, the correction terms $\beta_{m}$ are constructed from congruence data modulo $p^{\,k_{p,m}+1}$ (in particular via an inverse of $a$ modulo $p^{\,k_{p,m}+1}$; see the construction preceding [JKS21, Theorem 1.6]). Hence the reduced classes $\beta_{m}\bmod p^{\,k_{p,m}}$ are determined by $a\bmod p^{e+1}$, since $k_{p,m}\leq e$. Since $d_{p}(a)$ is determined by $a\bmod p^{e+1}$, the local $p$-shape is determined by $a\bmod p^{e+1}$.
Already among $p$-adic units one can choose
@@
with $p\nmid uu^{\prime}$, for which $r_{p}(a)=e-1$ while $r_{p}(a^{\prime})\geq e$, hence $d_{p}(a)\neq d_{p}(a^{\prime})$. Therefore the $p$-part of the shape changes, so $p^{e+1}$ is the least local modulus at $p$.
By the Chinese Remainder Theorem the global shape depends only on $a\bmod\prod_{p\mid n}p^{e_{p}+1}=n\cdot\mathrm{rad}(n)$, giving the optimal upper bound; local sharpness at every $p\mid n$ makes this modulus minimal.
@@
The point of view developed here—the local shape at each $p\mid n$ and its global gluing—yields several immediate applications. Firstly, we can show that one can compute, for any fixed $n$, a finite lookup table indexed by the classes $a\bmod M(n)$ recording the pp–denominator exponents and reduced corrections for every $p\mid n$. Thanks to Theorem 3.1, the table itself can be filled entirely locally using the explicit local recipes in [JKS21]. This offers a uniform and predictable alternative to ad hoc casework used in small degrees.
The second application concerns local tests for the $p$-part of the index $\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]]$ at primes $p\mid n$. When $p\parallel n$ and $v_{p}(a)=0$, the dichotomy between $a^{p-1}\not\equiv 1\pmod{p^{2}}$ and $a^{p-1}\equiv 1\pmod{p^{2}}$ decides whether $p$ divides $\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]]$ (equivalently, whether $\mathbb{Z}[\theta]$ is pp–maximal). More generally, for $p^{e}\parallel n$ and $v_{p}(a)=0$ the relevant invariant is $d_{p}(a)=\min\{\,v_{p}(a^{p-1}-1)-1,\,e\,\}$, and Theorem 3.1 shows that this (and with it all pp–denominator exponents and reduced corrections in an explicit basis) is determined by $a\bmod p^{e+1}$ and cannot be recovered from $a\bmod p^{e}$ alone.
@@
Historical remarks and related results
@@
When Dedekind’s criterion fails (i.e. some $p\mid[\mathcal{O}_{K}:\mathbb{Z}[\theta]]$), the modern local machine is based on Newton polygons at $p$, initiated by Ore in the 1920s and refined into a systematic algorithm by Montes and collaborators: one builds (higher-order) $p$-adic Newton polygons, reads residual polynomials along principal sides, and from this recovers $p$-adic factorization, indices, discriminants, and $p$-integral bases [ORE28, MN92, GMN12, GMN15]. This underlies much current algorithmic number theory around integral bases and prime decomposition. For practical $p$-integral basis extraction and index computations using Newton polygons, see ElFadil and his collaborators’ papers [EMN12, EL 20] and expositions cited therein. For the complexity and implementation discussions, take a look at [BNS13]. These provide worked implementations that agree with and conceptually support the $p$-local pieces used by Gaál–Remete and Jakhar–Khanduja–Sangwan.
@@
Section 2 introduces the shape of an integral basis and gives illustrative examples. Section 3 is preceded by our main local–to–global result (Theorem 3.1) and its direct consequences. We also include a tame relative version (Proposition 3.6) and examples. Section 4 develops applications: sharp local power-basis, explicit discriminant jumps, and counting/density results for prescribed shape classes, with worked quartic and sextic cases.
@@ Acknowledgments
We are grateful to the Morningside Center of Mathematics, Chinese Academy of Sciences, for its support and a stimulating research environment. We really appreciate the anonymous reviewer for their careful reading of the manuscript and for their valuable comments and suggestions. We thank István Gaál for his interest in our work and for his helpful comments.
Conflict of Interest
The author declares that there is no conflict of interest.
Data availability
This research is purely theoretical; no datasets were generated or analyzed. All mathematical results and worked examples are contained in the article.
@@ 2. Shape, periodicity, and minimal period
Throughout let $n\geq 3$ and $K_{a}=\mathbb{Q}(\theta)$ with $\theta^{n}=a$, where $a\in\mathbb{Z}$ is $n$th‑power‑free, $a\neq\pm 1$, and $x^{n}-a$ is irreducible over $\mathbb{Q}$. Write $n=\prod_{p}p^{e_{p}}$ and $\mathrm{rad}(n)=\prod_{p\mid n}p$. When we write an integral basis in the $\{1,\theta,\dots,\theta^{n-1}\}$–coordinates we mean a $\mathbb{Z}$–basis of the form
@@
Here $D_{m}$ denotes the (normalized) denominator of the $m$th basis element in these fixed coordinates, after clearing the factors supported on primes dividing $a$ into the numerator (cf. Remark 2.4), with $D_{m}\in\mathbb{Z}_{\geq 1}$, $\beta_{m}\in\mathbb{Z}[\theta]$ and linear in $\{1,\theta,\dots,\theta^{m-1}\}$. We use the following standard assumption in [JKS21] under which the explicit integral bases of $\mathcal{O}_{K_{a}}$ are stated. In particular, it holds whenever $a$ is squarefree or $(a,n)=1$.
@@
Write $n=\prod_{p}p^{e_{p}}$ and factor $a$ (uniquely up to sign) as
$a=\pm\prod_{j=1}^{n-1}a_{j}^{\,j},$
with the $a_{j}$ squarefree and pairwise coprime. For $0\leq m\leq n-1$ set
@@
For each $m$ and each $p\mid n$, let $k_{p,m}$ be the largest integer $k$ with $0\leq k\leq d_{p}(a)$ such that $m\geq n-n/p^{\,k}$ (put $k_{p,m}=0$ if no such $k$ exists). We obtain the following remarkable result, which is a reformulation of [JKS21, Theorem 1.6]. The correction terms $\beta_{m}$ (and hence their reductions modulo $p^{k_{p,m}}$) are given explicitly by the construction preceding [JKS21, Theorem 1.6], based on the auxiliary elements $\delta_{i,m}$ from [JKS21, Lemma 1.2]; see also [JKS21, Remark 1.7] for the freedom in choosing inverses.
@@
If $k_{p,m}=0$, then we regard the residue class $\beta_{m}\bmod p^{\,k_{p,m}}$ as trivial; in particular the pp–local shape depends only on those $\beta_{m}$ with $k_{p,m}>0$.
@@
Fix $n$ and let $N\geq 1$. We say that the shape is periodic modulo $N$ if, for all $n$th–power–free integers $a,a^{\prime}\in\mathbb{Z}$ for which $\mathbf{S}(a)$ and $\mathbf{S}(a^{\prime})$ are defined,
$a\equiv a^{\prime}\pmod{N}\qquad\Longrightarrow\qquad\mathbf{S}(a)=\mathbf{S}(a^{\prime}).$
A positive integer $N$ with this property is called a period (for fixed $n$). The minimal period is the least such $N$.
@@
Assume Hypothesis (H). Let $a$ be 4th‑power‑free. Put $r=v_{2}(a-1)-1$ when $a$ is odd, and r=1r=-1 when $2\mid a$. Then $d=\min\{r,2\}$. The general rule above gives
@@ 3. Main theorems
The proof of our main theorem uses only the explicit local invariants and bases in Jakhar–Khanduja–Sangwan [JKS21] (under Hypothesis (H)) and a Chinese remainder argument.
@@
Let $n\geq 3$, $a\in\mathbb{Z}$ be $n$th–power–free with $a\neq\pm 1$, and assume $x^{n}-a$ is irreducible over $\mathbb{Q}$, and $K_{a}=\mathbb{Q}(\theta)$ with $\theta^{n}=a$ satisfying hypothesis (H). For each prime $p\mid n$ write $p^{e_{p}}\parallel n$ and let $\mathbf{S}_{p}(a)$ and $\mathbf{S}(a)$ be the local and global shapes.
@@
    (Local determinacy.) For every $p\mid n$, the local shape $\mathbf{S}_{p}(a)$ depends only on the residue class of $a$ modulo $p^{\,e_{p}+1}$. Equivalently, if $a^{\prime}$ is another $n$th–power–free integer satisfying the same standing hypotheses (in particular (H) and irreducibility) and $a\equiv a^{\prime}\pmod{p^{\,e_{p}+1}}$, then $\mathbf{S}_{p}(a)=\mathbf{S}_{p}(a^{\prime})$.
@@
    and $M(n)$ is the least such modulus: if $N\geq 1$ has the property that for all $n$th–power–free $a,a^{\prime}$ satisfying the standing hypotheses,
    $a\equiv a^{\prime}\pmod{N}\qquad\Longrightarrow\qquad\mathbf{S}(a)=\mathbf{S}(a^{\prime}),$
    we get $M(n)\mid N$.